# Parameter-Efficient Fine-Tuning for Continual Learning: A Neural Tangent Kernel Perspective

### v2

@@ 6.4 Alternative Experiments
To rigorously evaluate and underscore the distinct advantages of the proposed components, a series of meticulously designed alternative experiments are conducted on the CIFAR100 dataset. Each experiment adheres to a stringent protocol, employing a fixed random seed of 0 to ensure reproducibility, while all experimental conditions and runtime environments are rigorously standardized to maintain consistency. A particular emphasis is placed on the incremental top-1 accuracy $A_{\tau}$ across tasks.
@@
The alternative experiments of sample size expansion methods encompasses two primary paradigms. Firstly, image-level augmentation techniques are explored, with Mixed-up [105], PuzzleMix [41], AutoAug [16], and RandAug [17] serving as representative methods. In this context, only the S2 component is retained, given its demonstrated superiority over S1 module on the CIFAR100 dataset, leveraging it to facilitate PEFT and the feature fusion from paired images. Despite achieving an expansion of the sample size at the image level, the results presented in Table VII reveal that this paradigm’s continual performance remains inferior to the network-level feature size expansion achieved through PEFT combinations. Furthermore, this paradigm incurs a higher computational cost, necessitating dual passes through both the backbone and subnetworks, as opposed to the single pass through the backbone and dual passes through the subnetworks required by the PEFT combinations. The second paradigm involves the amalgamation of distinct PEFT techniques, specifically IA3 [53], Compacter [58], and Side-Tuning [106]. As evidenced in Table VII, other PEFT combinations do not surpass the efficacy of ours. This disparity can be attributed to the fact that S1 module and S2 module extract information from the same image, albeit in the channel and spatial dimensions, respectively. This targeted extraction yields feature subspaces that are more discriminative and less redundant compared to those generated by other PEFT combinations that capture dataset-level biases, thereby contributing to superior performance.
In addition, systematic alternative experiments have been carried out concerning feature fusion, knowledge inheritance, orthogonality loss, and regularization loss. Comprehensive findings are summarized in Table VIII, Table IX, Table X, and Table XI respectively. Table VIII elucidates that the MSA technique, rigorously substantiated by theorem 4, stands out as the preeminent strategy for feature fusion. This method not only attains superior generalization loss but also excels in continual performance when juxtaposed with alternative fusion methods. Regarding knowledge inheritance, Table IX highlights the distinct superiority of the proposed Adaptive EMA mechanism, which facilitates effective retention and integration of historical knowledge, markedly outperforming its counterparts. The analysis presented in Table X reveals that, within the PEFT-CL scenario, task-level orthogonality suffices, diverging from conventional CL settings where class-level orthogonality is deemed essential. Lastly, the insights derived from Table XI affirm the validity of the adopted L2 regularization adjustment in solving saddle point dynamics in PEFT-CL, as delineated in Eq. 32. These findings collectively underscore the specificity and efficacy of our proposed components across various research dimensions.
@@ Appendix B NTK Dynamics in PEFT-CL
×$\left.\times(p_{\tau}-p_{\tau-1}^{*})-Y\right\|_{2}^{2}+\lambda\left\|p_{\tau}-p_{\tau-1}^{*}\right\|_{2}^{2}.$
@@
×$\quad\times(\phi_{\tau}(X)^{\top}\phi_{\tau}(X)+\lambda I)^{-1}(Y-f_{\tau-1}^{*}(X)),$
@@
×$\quad\times(\phi_{\tau}(X)^{\top}\phi_{\tau}(X)+\lambda I)^{-1}(Y-f_{\tau-1}^{*}(X)),$
@@
×$\quad\times(\Phi_{\tau}(X,X)+\lambda I)^{-1}(Y-f_{\tau-1}^{*}(X)).$
@@ Appendix C Task-Interplay Generalization in PEFT-CL
$f(x)=\left<\Phi(\cdot,x),f\right>_{\mathcal{H}}.$ (41)
@@
$||\tilde{f}_{\tau}^{*}||^{2}_{\mathcal{H}}$ $=\tilde{Y}_{\tau}^{\top}(\Phi_{\tau}(X,X)+\lambda I)^{-1}$ (45)
×$\quad\times\Phi_{\tau}(X,X)(\Phi_{\tau}(X,X)+\lambda I)^{-1}\tilde{Y}_{\tau},$
@@
×$\quad\times\Phi_{\tau}(X,X)(\Phi_{\tau}(X,X))^{-1}\tilde{Y}_{\tau},$
@@
$f_{T}(x)$ $=\sum\limits_{\tau=1}^{T}\sum\limits_{i=1}^{n_{\tau}}\alpha_{\tau}^{i}\left<\varphi_{\tau}(x),\varphi_{\tau}(x_{\tau}^{i})\right>_{\mathcal{H}},$ (46)
$=\sum\limits_{\tau=1}^{T}\sum\limits_{i=1}^{n_{\tau}}\alpha_{\tau}^{i}\left<\varphi_{\tau}(x_{\tau}^{i}),\varphi_{\tau}(x)\right>_{\mathcal{H}},$
$=\sum\limits_{\tau=1}^{T}\left<\sum\limits_{i=1}^{n_{\tau}}\alpha_{\tau}^{i}\varphi_{\tau}(x_{\tau}^{i}),\varphi_{\tau}(x)\right>_{\mathcal{H}},$
@@
$||\sum\limits_{i=1}^{n_{\tau}}\alpha_{\tau}^{i}\varphi_{\tau}(x_{\tau}^{i})||_{\mathcal{H}}^{2}$ $=\sum\limits_{i,j}\alpha_{\tau}^{i\top}\Phi_{\tau}(x_{\tau}^{i},x_{\tau}^{j})\alpha_{\tau}^{j}\leq G^{2}_{\tau},$ (47)
$\mathcal{F}_{T}\subset\{x\rightarrow\sum\limits_{\tau=1}^{T}\left<w_{\tau},\varphi_{\tau}(x)\right>_{\mathcal{H}},||w_{\tau}||_{\mathcal{H}}^{2}\leq G^{2}_{\tau}\}_{\mathcal{D}}:=\tilde{\mathcal{F}}_{T},$ (48)
Initially, the set ~$\tilde{\mathcal{F}}_{T}$ comprises functions characterized by the inner product between the feature mapping $\varphi_{\tau}(x)$ and the weight vector $w_{\tau}$ within the RKHS, in the form of $w_{\tau},\varphi_{\tau}(x)\right>_{\mathcal{H}}$. Accordingly, any arbitrary function $f_{\tau}^{*}(x)$ in $\mathcal{F}_{T}$ can be decomposed into the sum of the output of the previous task and the current task output change $f_{\tau}^{*}(x)=f_{\tau-1}^{*}(x)+\tilde{f}_{\tau}^{*}(x)$, and $\tilde{f}_{\tau}^{*}(x)$ is reconstructed into $w_{\tau},\varphi_{\tau}(x)\right>_{\mathcal{H}}$. Thus, as every function $f_{\tau}^{*}(x)$ in $\mathcal{F}_{T}$ can be reconstructed into the form found in ~$\tilde{\mathcal{F}}_{T}$, it can be concluded that $\mathcal{F}_{T}$ is a subset of ~$\tilde{\mathcal{F}}_{T}$.
Combining the computation method of Rademacher Complexity, we obtain the upper bound of $\hat{\mathcal{R}}(\tilde{\mathcal{F}})$,
$\hat{\mathcal{R}}(\mathcal{F})$ $=\mathbb{E}_{\epsilon}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f(x_{i})\right],$ (49)
^$\hat{\mathcal{R}}(\mathcal{F}_{T})$(~$\leq\hat{\mathcal{R}}(\tilde{\mathcal{F}}_{T}),$
$=\sum\limits_{\tau=1}^{T}\mathbb{E}_{\epsilon}\left[\sup_{||w_{\tau}||_{\mathcal{H}}^{2}\leq G_{\tau}^{2}}\left<w_{\tau},\frac{1}{n_{\tau}}\sum_{i=1}^{n_{\tau}}\epsilon_{i}\varphi_{\tau}(x_{\tau}^{i})\right>_{\mathcal{H}}\right],$ (50)
where $\epsilon_{i}$ are independently and identically distributed random variables, taking values of ±$\pm 1$. And since $\mathcal{F}_{T}$ is a subset of ~$\tilde{\mathcal{F}}_{T}$, its Rademacher Complexity is less than or equal to that of ~$\tilde{\mathcal{F}}_{T}$.
@@
^$\hat{\mathcal{R}}_{n}(\mathcal{F})$ $\leq\frac{2B}{n}\sqrt{\sum_{i=1}^{n}\mathbb{E}[k(X_{i},X_{i})]}.$ (52)
@@
[…]
@@
Clearly, the same argument applies with any independent, zero mean, unit variance random variables replacing the $g_{i}$, which gives the same bound for ^$\hat{\mathcal{R}}_{n}(\mathcal{F})$.
@@
^$\hat{\mathcal{R}}(\mathcal{F}_{T})$ $\leq\left[\sum\limits_{\tau=1}^{T}\frac{G_{\tau}}{n_{\tau}}\sqrt{\text{Tr}(\Phi_{\tau}(X,X))}\right]_{\mathcal{D}_{\tau}},$ (57)
@@
×$\quad\times\tilde{Y}_{\tau}-\tilde{Y}_{\tau}||_{2}^{2},$
@@
×$\quad\times(\Phi_{\tau}(X_{\tau},X_{\tau})+\lambda I)^{-1}\tilde{Y}_{\tau}-\tilde{Y}_{\tau}||_{2}^{2},$
@@
×$\quad\times\tilde{Y}_{\tau}-\tilde{Y}_{\tau}||_{2}^{2},$
@@
×$\quad\times\tilde{Y}_{\tau}||_{2}^{2},$
@@
$||\tilde{f}^{*}_{k}(X_{\tau})||_{2}^{2}=$ $\tilde{Y}_{k}^{\top}(\Phi_{k}(X_{k},X_{k})+\lambda I)^{-1}\Phi_{k}(X_{\tau},X_{k})$ (61)
×$\times\Phi_{k}(X_{\tau},X_{k})^{\top}(\Phi_{k}(X_{k},X_{k})+\lambda I)^{-1}\tilde{Y}_{k}.$
Then, the inequality for $||f_{T}^{*}(X_{\tau})-Y_{\tau}||_{2}^{2}$ is given by:
$\mathcal{L}_{S}(f_{T}^{*})$ $=||f_{T}^{*}(X_{\tau})-Y_{\tau}||_{2}^{2}$ (62)
@@
×$\quad\times\Phi_{k}(X_{\tau},X_{k})\Phi_{k}(X_{\tau},X_{k})^{\top}$
×$\quad\times(\Phi_{k}(X_{k},X_{k})+\lambda I)^{-1}\tilde{Y}_{k}\Big{]}.$
@@
Here, $\rho$ represents the Lipschitz constant. The term ^$\hat{\mathcal{R}}(\mathcal{F}_{T})$ refers to the empirical Rademacher complexity, as detailed in Eq. 57. $L_{S}(f^{*}_{T})$ represents the empirical loss in Eq. 62 and $\delta$ specifies the confidence level. While $c$ is a constant and $N$ denotes the total sample size.
@@ Appendix D Task-Intrinsic Generalization in PEFT-CL
×$\times\Lambda(\varphi(x)\varphi(x)^{\top}+\lambda I)^{-1}w^{*}\Big{>}_{x\in D_{\tau}}.$
@@
×$\quad\left.\times\Lambda(\varphi(x)\varphi(x)^{\top}+\lambda I)^{-1}w^{*}\right>_{x\in D_{\tau}},$
=$\Lambda^{-\frac{1}{2}}w^{*}w^{*\top}\Lambda^{-\frac{1}{2}}\right>_{x\in D_{\tau}}$
×$(\lambda\Lambda^{\frac{1}{2}}(\varphi(x)\varphi(x)^{\top}+\lambda I)^{-1}\Lambda^{\frac{1}{2}})^{2}\right>_{x\in D_{\tau}},$
=$\Lambda^{-\frac{1}{2}}w^{*}w^{*\top}\Lambda^{-\frac{1}{2}}\right>_{x\in D_{\tau}}$
×$((\frac{1}{\lambda}O(x)O(x)^{\top}+\Lambda^{-1})^{-1})^{2}\right>_{x\in D_{\tau}},$
@@
$U(s+1,z)\right>_{x\in D_{\tau}}=\left<\left(U(s,z)^{-1}+\frac{1}{\lambda}O(x)O(x)^{\top}\right)^{-1}\right>_{x\in D_{\tau}},$ (73)
[…]
+$(\lambda I+O(x)^{\top}U(s,z)O(x))^{-1}O(x)^{\top}U(s,z)\right>_{x\in D_{\tau}},$
=$U(s,z)\right>_{x\in D_{\tau}}-\left<\frac{U(s,z)O(x)O(x)^{\top}U(s,z)}{\lambda+O(x)^{\top}U(s,z)O(x)}\right>_{x\in D_{\tau}},$
@@
[…]
@@
×$\times\left(1-\frac{s}{(\lambda+TU(s,0))^{2}}\frac{\partial TU(s,0)}{\partial z}\right).$
@@
×$\quad\times\left(1-\frac{s}{(\lambda+TU(s,0))^{2}}\frac{\partial TU(s,0)}{\partial z}\right).$
@@
×$\times(1-\frac{ms}{(\lambda+TU(s,0))^{2}})^{-1}.$
@@
×$\quad\times(1-\frac{ms}{(\lambda+TU(s,0))^{2}})^{-1},$
@@
×$\quad\times(1-\frac{ms}{(\lambda+TU(s))^{2}})^{-1}.$

### v3

@@ 6.4 Alternative Experiments
To rigorously evaluate and underscore the distinct advantages of the proposed components, a series of meticulously designed alternative experiments are conducted on the CIFAR100 dataset. Each experiment adheres to a stringent protocol, employing the fixed seeds of 0-4 to ensure reproducibility, while all experimental conditions and runtime environments are rigorously standardized to maintain consistency. A particular emphasis is placed on the incremental top-1 accuracy $A_{\tau}$ across tasks.
@@
The alternative experiments of sample size expansion methods encompasses two primary paradigms. Firstly, image-level augmentation techniques are explored, with Mixed-up [105], PuzzleMix [41], AutoAug [16], and RandAug [17] serving as representative methods. In this context, only the S2 component is retained, given its demonstrated superiority over S1 module, leveraging it to facilitate PEFT and the feature fusion from paired images. Despite achieving an expansion of the sample size at the image level, the results presented in Table VII reveal that this paradigm’s continual performance remains inferior to the network-level feature size expansion achieved through PEFT combinations. The second paradigm involves the amalgamation of distinct PEFT techniques, specifically IA3 [53], Compacter [58], and Side-Tuning [106]. As evidenced in Table VII, other PEFT combinations do not surpass the efficacy of ours. This disparity can be attributed to the fact that S1 module and S2 module extract information from the same image, albeit in the channel and spatial dimensions, respectively. This targeted extraction yields feature subspaces that are more discriminative and less redundant compared to those generated by other PEFT combinations that capture dataset-level biases, thereby contributing to superior performance.
In addition, we conduct systematic comparative studies on feature fusion strategies, knowledge inheritance mechanisms, orthogonality constraints, and regularization schemes under the proposed dual-stream architecture with a shared and frozen backbone. These experiments are designed to evaluate the effectiveness of individual components independently of the backbone execution flow, and all results are obtained under identical training and inference settings. The comparative results are summarized in Table VIII, Table IX, Table X, and Table XI, respectively. Table VIII shows that the proposed MSA-based fusion strategy achieves superior continual performance compared to alternative fusion designs, which is consistent with the generalization analysis presented in theorem 4. Regarding knowledge inheritance, the results in Table IX demonstrate the effectiveness of the proposed Adaptive EMA mechanism in preserving and integrating historical knowledge across tasks. Table X further indicates that, in the PEFT-CL setting, enforcing task-level orthogonality is sufficient, differing from conventional CL scenarios where class-level orthogonality is often required. Finally, Table XI validates the adopted L2 regularization scheme in stabilizing the optimization dynamics of PEFT-CL, as characterized in Eq. 32. Collectively, these results confirm that the proposed design choices are well-suited to the dual-stream continual learning framework and contribute consistently to performance improvements.
@@ Appendix B NTK Dynamics in PEFT-CL
$\left.\times(p_{\tau}-p_{\tau-1}^{*})-Y\right\|_{2}^{2}+\lambda\left\|p_{\tau}-p_{\tau-1}^{*}\right\|_{2}^{2}.$
@@
$\qquad\times(\phi_{\tau}(X)^{\top}\phi_{\tau}(X)+\lambda I)^{-1}(Y-f_{\tau-1}^{*}(X)),$
@@
$\qquad\times(\phi_{\tau}(X)^{\top}\phi_{\tau}(X)+\lambda I)^{-1}(Y-f_{\tau-1}^{*}(X)),$
@@
$\qquad\times(\Phi_{\tau}(X,X)+\lambda I)^{-1}(Y-f_{\tau-1}^{*}(X)).$
@@ Appendix C Task-Interplay Generalization in PEFT-CL
$f(x)=\left<\Phi(\cdot,x),f\right$$_{\mathcal{H}}$. (41)
@@
[…]
$\quad\times\Phi_{\tau}(X,X)(\Phi_{\tau}(X,X)+\lambda I)^{-1}\tilde{Y}_{\tau},$
@@
$\quad\times\Phi_{\tau}(X,X)(\Phi_{\tau}(X,X))^{-1}\tilde{Y}_{\tau},$
@@
$f_{T}(x)$ $=\sum\limits_{\tau=1}^{T}\sum\limits_{i=1}^{n_{\tau}}\alpha_{\tau}^{i}\left<\varphi_{\tau}(x),\varphi_{\tau}(x_{\tau}^{i})\right$$_{\mathcal{H}}$, (46)
$=\sum\limits_{\tau=1}^{T}\sum\limits_{i=1}^{n_{\tau}}\alpha_{\tau}^{i}\left<\varphi_{\tau}(x_{\tau}^{i}),\varphi_{\tau}(x)\right$$_{\mathcal{H}}$,
$=\sum\limits_{\tau=1}^{T}\left<\sum\limits_{i=1}^{n_{\tau}}\alpha_{\tau}^{i}\varphi_{\tau}(x_{\tau}^{i}),\varphi_{\tau}(x)\right$$_{\mathcal{H}}$,
@@
[…]
Initially, the set $\tilde{\mathcal{F}}_{T}$ comprises functions characterized by the inner product between the feature mapping $\varphi_{\tau}(x)$ and the weight vector $w_{\tau}$ within the RKHS, in the form of $\left<w_{\tau},\varphi_{\tau}(x)\right$$_{\mathcal{H}}$. Accordingly, any arbitrary function $f_{\tau}^{*}(x)$ in $\mathcal{F}_{T}$ can be decomposed into the sum of the output of the previous task and the current task output change $f_{\tau}^{*}(x)=f_{\tau-1}^{*}(x)+\tilde{f}_{\tau}^{*}(x)$, and $\tilde{f}_{\tau}^{*}(x)$ is reconstructed into $\left<w_{\tau},\varphi_{\tau}(x)\right$$_{\mathcal{H}}$. Thus, as every function $f_{\tau}^{*}(x)$ in $\mathcal{F}_{T}$ can be reconstructed into the form found in $\tilde{\mathcal{F}}_{T}$, it can be concluded that $\mathcal{F}_{T}$ is a subset of $\tilde{\mathcal{F}}_{T}$.
[…]
$\hat{\mathcal{R}}(\mathcal{F}_{T})$ $\leq\hat{\mathcal{R}}(\tilde{\mathcal{F}}_{T}),$
[…]
where $\epsilon_{i}$ are independently and identically distributed random variables, taking values of $\pm 1$. And since $\mathcal{F}_{T}$ is a subset of $\tilde{\mathcal{F}}_{T}$, its Rademacher Complexity is less than or equal to that of $\tilde{\mathcal{F}}_{T}$.
@@
$\hat{\mathcal{R}}_{n}(\mathcal{F})$ $\leq\frac{2B}{n}\sqrt{\sum_{i=1}^{n}\mathbb{E}[k(X_{i},X_{i})]}.$ (52)
@@
Suppose that $\mathcal{H}$ is a Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and induced norm $\|\cdot\|$, and the kernel $k$ has feature map $\phi:\mathcal{X}\to\mathcal{H}$. Let $g_{1},\ldots,g_{n}$ be independent standard normal random variables. Then
@@
Clearly, the same argument applies with any independent, zero mean, unit variance random variables replacing the $g_{i}$, which gives the same bound for $\hat{\mathcal{R}}_{n}(\mathcal{F})$.
@@
$\hat{\mathcal{R}}(\mathcal{F}_{T})$ $\leq\left[\sum\limits_{\tau=1}^{T}\frac{G_{\tau}}{n_{\tau}}\sqrt{\text{Tr}(\Phi_{\tau}(X,X))}\right]_{\mathcal{D}_{\tau}},$ (57)
@@
$\qquad\times\tilde{Y}_{\tau}-\tilde{Y}_{\tau}||_{2}^{2},$
@@
$\qquad\times(\Phi_{\tau}(X_{\tau},X_{\tau})+\lambda I)^{-1}\tilde{Y}_{\tau}-\tilde{Y}_{\tau}||_{2}^{2},$
@@
$\qquad\times\tilde{Y}_{\tau}-\tilde{Y}_{\tau}||_{2}^{2},$
@@
$\qquad\times\tilde{Y}_{\tau}||_{2}^{2},$
@@
[…]
$\times\Phi_{k}(X_{\tau},X_{k})^{\top}(\Phi_{k}(X_{k},X_{k})+\lambda I)^{-1}\tilde{Y}_{k}.$
Then, the inequality for $f_{T}^{*}(X_{\tau})-Y_{\tau}||_{2}^{2}$ is given by:
[…]
@@
$\quad\times\Phi_{k}(X_{\tau},X_{k})\Phi_{k}(X_{\tau},X_{k})^{\top}$
$\quad\times(\Phi_{k}(X_{k},X_{k})+\lambda I)^{-1}\tilde{Y}_{k}\Big].$
@@
Here, $\rho$ represents the Lipschitz constant. The term $\hat{\mathcal{R}}(\mathcal{F}_{T})$ refers to the empirical Rademacher complexity, as detailed in Eq. 57. $L_{S}(f^{*}_{T})$ represents the empirical loss in Eq. 62 and $\delta$ specifies the confidence level. While $c$ is a constant and $N$ denotes the total sample size.
@@ Appendix D Task-Intrinsic Generalization in PEFT-CL
$\times\Lambda(\varphi(x)\varphi(x)^{\top}+\lambda I)^{-1}w^{*}\Big>_{x\in D_{\tau}}.$
@@
$\qquad\left.\times\Lambda(\varphi(x)\varphi(x)^{\top}+\lambda I)^{-1}w^{*}\right>_{x\in D_{\tau}},$
$=\left<\Lambda^{-\frac{1}{2}}w^{*}w^{*\top}\Lambda^{-\frac{1}{2}}\right>_{x\in D_{\tau}}$
$\qquad\times\left<(\lambda\Lambda^{\frac{1}{2}}(\varphi(x)\varphi(x)^{\top}+\lambda I)^{-1}\Lambda^{\frac{1}{2}})^{2}\right>_{x\in D_{\tau}},$
$=\left<\Lambda^{-\frac{1}{2}}w^{*}w^{*\top}\Lambda^{-\frac{1}{2}}\right>_{x\in D_{\tau}}$
$\qquad\times\left<((\frac{1}{\lambda}O(x)O(x)^{\top}+\Lambda^{-1})^{-1})^{2}\right>_{x\in D_{\tau}},$
@@
$\left<U(s+1,z)\right>_{x\in D_{\tau}}=\left<\left(U(s,z)^{-1}+\frac{1}{\lambda}O(x)O(x)^{\top}\right)^{-1}\right>_{x\in D_{\tau}},$ (73)
$=\left<U(s,z)\right>_{x\in D_{\tau}}-\left<U(s,z)O(x)\right>_{x\in D_{\tau}}$
$\quad+\left<(\lambda I+O(x)^{\top}U(s,z)O(x))^{-1}O(x)^{\top}U(s,z)\right>_{x\in D_{\tau}},$
$=\left<U(s,z)\right>_{x\in D_{\tau}}-\left<\frac{U(s,z)O(x)O(x)^{\top}U(s,z)}{\lambda+O(x)^{\top}U(s,z)O(x)}\right>_{x\in D_{\tau}},$
@@
$\left<U(s+1,z)\right>_{x\in D_{\tau}}\approx\left<U(s,z)\right>_{x\in D_{\tau}}-\frac{\left<U(s,z)^{2}\right>_{x\in D_{\tau}}}{\lambda+\text{Tr}\left<U(s,z)\right>_{x\in D_{\tau}}}.$ (74)
@@
$\times\left(1-\frac{s}{(\lambda+TU(s,0))^{2}}\frac{\partial TU(s,0)}{\partial z}\right).$
@@
$\quad\times\left(1-\frac{s}{(\lambda+TU(s,0))^{2}}\frac{\partial TU(s,0)}{\partial z}\right).$
@@
$\times(1-\frac{ms}{(\lambda+TU(s,0))^{2}})^{-1}.$
@@
$\quad\times(1-\frac{ms}{(\lambda+TU(s,0))^{2}})^{-1},$
@@
$\quad\times(1-\frac{ms}{(\lambda+TU(s))^{2}})^{-1}.$